AcornTheNut:

remind me, why can't you divide 0 (nothing) by a number with a sort of value?

Aspie7:

anything divided by 0 is undefined

ChloeHamilton:

it's like multiplying by zero. if you multiply something by zero, you get zero. division is basically the opposite of multiplication. hence, if you tried to divide something by zero, there isn't a known answer rather than zero

axie:

So dividing is kinda like putting x amount of things in x amount of groups equally. So imagine taking a bag of m&ms and splitting it evenly with 4 friends. If u didn't have any friends then it's impossible and if you dont have m&ms then it's also impossible

axie:

It would look like 500/5

axie:

500/4****

59mandatory:

because anything times 0 is 0.

surjithayer:

let 0/a=b, \[a \neq 0\] \[0=a \times b\] if product of two numbers is zero then either one or both a and b is zero. \[a \neq 0,so ~b = 0\]

poopoopeepee:

as surj said though, you can... 0/(anything not 0) = 0 Unless you meant the other way around, in which case consider this: (anything including 0)/0 = undefined The first case you can multiply as surj as shown, so it's 0 The second case must be undefined, because if you multiply: \(a/0=b\) \(a=b*0\). If a is 0, b could literally be anything, so it's undefined if a isn't 0, b can't be anything, so it's still undefined.

AcornTheNut:

thank you everyone